OpenStudy (anonymous):

what is the limit of h/tan(h) as h approches 0?

hartnn (hartnn):

write tan h as sin x / cos x

OpenStudy (anonymous):

\[\frac{ h }{ \frac{ \sinh }{ \cosh } }\]

hartnn (hartnn):

write that as cos x /(sin x/x) did u understand how ?

OpenStudy (anonymous):

wouldn't it be \[\frac{ hcos(h) }{ \sin(h) }\]

hartnn (hartnn):

oops, sorry i used x instead. yes, and u can bring that h in denominator of sin h to get cos h/(sin h/h) ok ?

OpenStudy (anonymous):

how could I bring the h to the denominator?

hartnn (hartnn):

divide h in numerator and denominator

OpenStudy (anonymous):

the answer is 1. how would i get that?

hartnn (hartnn):

put h=0 in cos h (cos 0 = ?) use the standard limit sin x/x = 1 for denominator

OpenStudy (anonymous):

ok thanks

hartnn (hartnn):

welcome ^_^ i hope u understand all steps.